# Alleviating Exposure Bias in Diffusion Models through Sampling with Shifted Time Steps

### v5

@@ 1 Introduction
  • We propose a new sampling method called Time-Shift Sampler to alleviate the exposure bias problem, which avoids retraining the models. Our method can be seamlessly integrated to existing sampling method by only introducing minimal computational cost.
@@ 2.1 Background: Diffusion Probabilistic Models
where $\tilde{\beta}_{t}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t}$ and $\tilde{\mu}_{t}=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}x_{0}+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}x_{t}$. Considering Equation 2, $\tilde{\mu}_{t}$ can be further reformulated as $\tilde{\mu}_{t}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{t})$. During training, a time-dependent neural network is optimized by either learning $\tilde{\mu}_{t}$ or $\epsilon_{t}$. Empirically, Ho et al. (2020) observes that predicting $\epsilon_{t}$ works better. The learning target is to optimize the variational lower bound of the negative log-likelihood, which could also be interpreted as minimizing the KL divergence between the forward and backward process. In practice, Ho et al. (2020) further simplifies the loss function as:
@@ 2.2 The Exposure Bias in Diffusion Probabilistic Models
In the second experiment, given a specific number of sampling steps, we compute the mean squared errors between the predicted samples and the ground truth samples at each step, as presented in Figure 3. Details for plotting this figure are presented in Appendix B. It can be seen that the evolution of prediction errors adheres to a consistent pattern: initially decreasing before incrementally accumulating as the sampling process progresses. This phenomenon may be attributed to two possible factors: (1) The sampling process originates from white noise, which does not contain any information about the resultant sample distributions. In early stages, with fewer sampling steps, the error accumulation is less serious, thus the network gradually shapes the predicted distribution into the target distribution. (2) In the later stages, the network is more robust to the noisy inputs as discussed above. However, due to the exposure bias, the network inevitably makes errors at each step and this errors accumulate along the sequence, resulting in a slow but steady progress in error accumulation and larger errors in the end.
In conclusion, the above two experiments together show that during backward sampling process, the accumulated prediction error, which arises from the exposure bias and the capability of the network, could strongly impact the final results. This demonstrates the importance of alleviating exposure bias in DPM, which could potentially lead to improved results.
@@ 3 Alleviating Exposure Bias via Time Step Shifting
Let $\hat{x}_{t}$ represent a given state and $\hat{x}_{t-1}$ represent the predicted subsequent state. We assume $t-1$ is sufficiently large such that the distribution of $\hat{x}_{t-1}$ is still close to the initialized normal distribution with a diagonal covariance matrix. In addition, the selected time step $t_{s}$ to couple with $\hat{x}_{t-1}$ is among those time steps closely surrounding $t-1$  [^1] If $t_{s}$ is close to $t-1$, then $\sqrt{\bar{\alpha}_{t-1}}-\sqrt{\bar{\alpha}_{t_{s}}}\approx 0$. See details in Appendix J.1. Then the optimal $t_{s}$ should have the following variance:
@@
The derivation of Theorem 3.1 mainly follows two steps: Firstly, we optimize the KL divergence between $x_{t_{s}}$ and $\hat{x}_{t-1}$ to obtain the variance of $x_{t_{s}}$. Secondly, we establish the relationship between the variance within a single sample of $\hat{x}_{t-1}$ and the variance of $\hat{x}_{t-1}$ [^2] The latter refers to the variance computed by considering corresponding elements across samples of $\hat{x}_{t-1}$.. The results articulated in Theorem 3.1 could be further simplified to $\sigma_{t_{s}}\approx\sigma_{t-1}$, when $t$ is large and given the assumption that the network prediction error at the current time step is minimal. This assumption has been found to hold well in practice.
@@
Based on the findings of Theorem 3.1, we propose the Time-Shift Sampler, a method that can be seamlessly incorporated into existing sampling algorithms, such as Denoising Diffusion Implicit Models (DDIM) (Song et al., 2021), Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) or sampling methods based on high-order numerical solvers (Liu et al., 2022a). Moreover, in light of the earlier discussion that the model’s predictions of nearby time steps tend to converge to the same distribution during the later stage of the inference process and the condition of large $t$ in the derivation in Appendix J.1, we remove the time-shift operation when the time step is smaller then a predefined cutoff time step [^3] We discuss the influence of the window size and the cutoff value in Sec. 5.4.. Our algorithm is detailed in Algorithm 3. Specifically, given a trained DPM $\epsilon_{\theta}$, we sample for $x_{t},t=1,2,\ldots,T$ using arbitrarily any sampling method. For each time step $t>t_{c}$, where $t_{c}$ is the cutoff threshold, we replace the next time step $t_{next}$ with the time step $t_{s}$ that best couples with the variance of $x_{t-1}$ within a window of size $w$. In the search of such $t_{s}$, we first take a sampling step with the current $t_{next}$ to get $x_{t-1}$, which is used to compute the variance of $x_{t-1}$ as $var(x_{t-1})$. Then we get the variance $\Sigma$ of the time steps within the window. The optimal $t_{s}$ can be obtained via $\arg\min_{\tau}||var(x_{t-1}),\sigma_{\tau}||$, for $\sigma_{\tau}\in\Sigma$. Finally, the obtained $t_{s}$ is passed to the next sampling iteration as $t_{next}$. We repeat this process until $t<t_{c}$, after which we perform the conventional sampling steps from the sampling method of choice.
@@ 5.3 Comparison with Training-Required Method ADM-IP
Table 2: Performance comparison on CIFAR-10 with ADM and ADM-IP as the backbone models.
@@ 6 Related Work
Denoising Diffusion Probabilistic Model. The denoising diffusion probabilistic model (DDPM) was first introduced by Sohl-Dickstein et al. (2015) and further advanced by Nichol & Dhariwal (2021), where they include variance learning in the model and optimize it with a new weighted variational bound. Song et al. (2020) further connect DDPMs with stochastic differential equations by considering DDPMs with infinitesimal timesteps. They also find that both score-based generative models (Song & Ermon, 2019) and that DDPMs can be formulated by stochastic differential equations with different discretization. Some variants of DDPMs are introduced recently, including the variational diffusion model (VDM) Kingma et al. (2021) that also learns the forward process, and consistency model (Song et al., 2023) and rectified flow model (Liu et al., 2022b), both of which aim to learn to generate natural images in one inference step.
@@ Appendix B Figure Details
Figure 3 shows the mean square error (MSE) between the prediction and ground truth at each step in the backward process. Given an image denoted as $x_{0}$, by applying Equation 2 we could obtain a sequence of $x_{t},t=1,2,\cdots,T-1$. Taking each $x_{t}$ and the paired time step $t$ to run the backward process, we would obtain a sequence of predicted $\hat{x}_{0}^{t},t=1,2,\cdots,T-1$. Ideally, we would expect all these predicted $\hat{x}_{0}^{t},t=1,2,\cdots,T-1$, to be exactly equal to the ground truth $x_{0}$, as they are generated using the given $x_{0}$. However, this is not the case in practice. In our experiments, we found that only when $t<t_{s}$ (around 650 steps in our experiment using DDIM) we could obtain the original $x_{0}$ by running the backward process with paired $(x_{t},t)$. For $t>t_{s}$, the image created using $(x_{t},t)$ differs from the original $x_{0}$. This observation also reveals that the image generation process of diffusion models basically contains two stages. In the first stage, the model moves the Gaussian distribution towards the image distribution and no modes are presented at this stage, which means we can not know which images will be generated. In the second stage, the prediction shows clear patterns and modes are presented. We can predict which images will be generated following the backward process. This observation led us to divide the error computation into two stages. The full explanation, including the equations, is shown in the Figure 8
@@ Appendix C Additional Experimental Setup
Following DDIM (Song et al., 2021), we use two types of time step selection procedures, namely uniform and quadratic. For comparing different sampling methods, the architecture of our models follows the one of DDPM (Ho et al., 2020). For all datasets, we use the same pretrained models for evaluating the sampling methods. For CIFAR-10 and LSUN-bedroom, we obtain the checkpoints from the original DDPM implementation; for CelebA, we adopt the DDIM checkpoints. To carry out the comparison between ADM (Dhariwal & Nichol, 2021), ADM-IP Ning et al. (2023) and our method, we choose the ADM architecture and the checkpoints from ADM-IP.
@@ Appendix D Error Analysis
We provide here more examples of prediction errors on training samples for different sampling steps for the different datasets. Figure 9 presents the prediction errors obtained in the CelebA dataset, which show similar pattern as those of the CIFAR-10 dataset as dipicted in Figure 3.
@@ Appendix F Sampling Time Comparison
We report the detailed comparison on sampling time using different sampling methods in Table 3 and 4
@@ Appendix G Case Study
The proposed novel sampling method is motivated by the theoretical and empirical analysis of the exposure bias problem in DDPM. Experimental results show that our sampling method can effectively improve the quality of the generated images compared to images generated with the original DDIM/DDPM models quantatively measured with the FID score. We present the comparison of the generation chain of TS-DDIM (our method) and DDIM in Figure 11. It can be seen that TS-DDIM shifted most of the time steps before reaching the cutoff value, which yields much better generation quality of the image of a horse. Additional qualitative examples will be presented in the Appendix I.
@@
Figure 12: Comparison of timestep trajectory for TS-DDIM and DDIM on CIFAR-10 with DDPM as backbone using 10 sampling steps and uniform time selection. We use the red dashed line and blue line to present the time step trajectories for the original DDIM and our TS-DDIM, respectively. To improve visibility, we also zoom in for time step between 300 to 600.
@@ Appendix H Qualitative Examples
In this section we present the example images generated using different sampling method with uniform time selection procedure for varying sampling time steps. Generated examples can be found in Figure 13,  14,  15,  16,  17 and  18. It can be seen that we can generate images with good quality for less than 10 sampling steps.
@@ Appendix I Additional Results
We present more results on LSUN-bedroom (Yu et al., 2016) in Table 5. Limited by computational resources, we do not properly tune the parameters, i.e., window size and cutoff values, for the LSUN-bedroom. We leave the exploration of a more efficient tuning strategy on high-resolution images for future work.
@@
Table 6: Comparison with Analytic-DPM on CIFAR-10 (32×32) and CelebA (64×64)
@@
We further integrate our method into the DPM-solver (Lu et al., 2022) and DEIS (Zhang & Chen, 2023) to showcase its versatility. The results are presented in Table 9. Our model can still improve the performance of both DPM-solver and DEIS samplers. However, in comparison to DDIM, our method again provides substantial improvements although the improvements for these samplers are a bit smaller than for the other tested samplers.This could be attributed to the fact that both the DPM-solver and DEIS utilize the particular structure of the semi-linear ODE, which already largely reduces the error of estimating $x_{t}$.
@@
Finally, we report the performance of our method on text-to-image generation on MSCOCO val2017 in Table 10
@@
In this section, we prove the Theorem 3.1.
@@
$=\frac{1}{2}(d\log(1-\overline{\alpha}_{t_{s}})+\frac{d}{1-\overline{\alpha}_{t_{s}}}Tr(\hat{\Sigma}_{t-1})+\frac{1}{1-\overline{\alpha}_{t_{s}}}||\hat{\mu}_{t-1}-\mu_{t_{s}}||^{2})+C$
@@
$=\frac{1}{2}(d\log(1-\overline{\alpha}_{t_{s}})+\frac{d}{1-\overline{\alpha}_{t_{s}}}Tr(\hat{\Sigma}_{t-1})+\frac{1}{1-\overline{\alpha}_{t_{s}}}||(\sqrt{\overline{\alpha}_{t-1}}-\sqrt{\overline{\alpha}_{t_{s}}})x_{0}+e||^{2})+C$
@@
$\mathcal{D}_{KL}(p(\hat{x}_{t-1}|\hat{x}_{t})||q(x_{t_{s}}|x_{0}))\approx\frac{1}{2}(d\log(1-\overline{\alpha}_{t_{s}})+\frac{d}{1-\overline{\alpha}_{t_{s}}}Tr(\hat{\Sigma}_{t-1})+\frac{1}{1-\overline{\alpha}_{t_{s}}}||e||^{2})+C$ (12)
@@
$\Sigma_{t_{s}}=Tr(\hat{\Sigma}_{t-1})+\frac{1}{d}||e||^{2}$ (13)
@@
In the above derivation, we assume that when $t$ is large we can reach Equation 21. This assumption corresponds to the cutoff mechanism in our proposed algorithm, where we stop conducting time shift when $t$ is small, as the assumption does not hold and we are not able to estimate the $\hat{\Sigma}_{\hat{x}_{t-1}}$ in Equation 13.

### v8

@@ 1 Introduction
  • We propose a new sampling method called Time-Shift Sampler to alleviate the exposure bias problem, which avoids retraining the models. Our method can be seamlessly integrated into existing sampling methods by only introducing minimal computational cost.
@@ 2.1 Background: Diffusion Probabilistic Models
where $\tilde{\beta}_{t}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t}$ and $\tilde{\mu}_{t}=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}x_{0}+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}x_{t}$. Considering Equation 2, $\tilde{\mu}_{t}$ can be further reformulated as $\tilde{\mu}_{t}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{t})$. During training, a time-dependent neural network is optimized by either learning $\tilde{\mu}_{t}$ or $\epsilon_{t}$. Empirically, Ho et al. (2020) observe that predicting $\epsilon_{t}$ works better. The learning target is to optimize the variational lower bound of the negative log-likelihood, which could also be interpreted as minimizing the KL divergence between the forward and backward process. In practice, Ho et al. (2020) further simplify the loss function as:
@@ 2.2 The Exposure Bias in Diffusion Probabilistic Models
In the second experiment, given a specific number of sampling steps, we compute the mean squared errors between the predicted samples and the ground truth samples at each step, as presented in Figure 3. Details for plotting this figure are presented in Appendix B. It can be seen that the evolution of prediction errors adheres to a consistent pattern: initially decreasing before incrementally accumulating as the sampling process progresses. This phenomenon may be attributed to two possible factors: (1) The sampling process originates from white noise, which does not contain any information about the resultant sample distributions. In early stages, with fewer sampling steps, the error accumulation is less serious, thus the network gradually shapes the predicted distribution into the target distribution. (2) In the later stages, the network is more robust to the noisy inputs as discussed above. However, due to the exposure bias, the network inevitably makes errors at each step and these errors accumulate along the sequence, resulting in a slow but steady progress in error accumulation and larger errors in the end.
In conclusion, the above two experiments together show that during the backward sampling process, the accumulated prediction error, which arises from the exposure bias and the capability of the network, could strongly impact the final results. This demonstrates the importance of alleviating exposure bias in DPM, which could potentially lead to improved results.
@@ 3 Alleviating Exposure Bias via Time Step Shifting
Let $\hat{x}_{t}$ represent a given state and $\hat{x}_{t-1}$ represent the predicted subsequent state. We assume $t-1$ is sufficiently large such that the distribution of $\hat{x}_{t-1}$ is still close to the initialized normal distribution with a diagonal covariance matrix. In addition, the selected time step $t_{s}$ to couple with $\hat{x}_{t-1}$ is among those time steps closely surrounding $t-1$.  [^1] If $t_{s}$ is close to $t-1$, then $\sqrt{\bar{\alpha}_{t-1}}-\sqrt{\bar{\alpha}_{t_{s}}}\approx 0$. See details in Appendix J.1 Then the optimal $t_{s}$ should have the following variance:
@@
The derivation of Theorem 3.1 mainly follows two steps: Firstly, we optimize the KL divergence between $x_{t_{s}}$ and $\hat{x}_{t-1}$ to obtain the variance of $x_{t_{s}}$. Secondly, we establish the relationship between the variance within a single sample of $\hat{x}_{t-1}$ and the variance of $\hat{x}_{t-1}$.  [^2] The latter refers to the variance computed by considering corresponding elements across samples of $\hat{x}_{t-1}$. The results articulated in Theorem 3.1 could be further simplified to $\sigma_{t_{s}}\approx\sigma_{t-1}$, when $t$ is large and given the assumption that the network prediction error at the current time step is minimal. This assumption has been found to hold well in practice.
@@
Based on the findings of Theorem 3.1, we propose the Time-Shift Sampler, a method that can be seamlessly incorporated into existing sampling algorithms, such as Denoising Diffusion Implicit Models (DDIM) (Song et al., 2021), Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) or sampling methods based on high-order numerical solvers (Liu et al., 2022a). Moreover, in light of the earlier discussion that the model’s predictions of nearby time steps tend to converge to the same distribution during the later stage of the inference process and the condition of large $t$ in the derivation in Appendix J.1, we remove the time-shift operation when the time step is smaller then a predefined cutoff time step. [^3] We discuss the influence of the window size and the cutoff value in Sec. 5.4. Our algorithm is detailed in Algorithm 3. Specifically, given a trained DPM $\epsilon_{\theta}$, we sample for $x_{t},t=1,2,\ldots,T$ using arbitrarily any sampling method. For each time step $t>t_{c}$, where $t_{c}$ is the cutoff threshold, we replace the next time step $t_{next}$ with the time step $t_{s}$ that best couples with the variance of $x_{t-1}$ within a window of size $w$. In the search of such $t_{s}$, we first take a sampling step with the current $t_{next}$ to get $x_{t-1}$, which is used to compute the variance of $x_{t-1}$ as $var(x_{t-1})$. Then we get the variance $\Sigma$ of the time steps within the window. The optimal $t_{s}$ can be obtained via $\arg\min_{\tau}||var(x_{t-1}),\sigma_{\tau}||$, for $\sigma_{\tau}\in\Sigma$. Finally, the obtained $t_{s}$ is passed to the next sampling iteration as $t_{next}$. We repeat this process until $t<t_{c}$, after which we perform the conventional sampling steps from the sampling method of choice.
@@ 5.3 Comparison with Training-Required Method ADM-IP
Table 2: Performance comparison on CIFAR-10 with ADM and ADM-IP as backbone models.
@@ 6 Related Work
Denoising Diffusion Probabilistic Model. The denoising diffusion probabilistic model (DDPM) was first introduced by Sohl-Dickstein et al. (2015) and further advanced by Nichol & Dhariwal (2021), where they include variance learning in the model and optimize it with a new weighted variational bound. Song et al. (2020) further connect DDPMs with stochastic differential equations by considering DDPMs with infinitesimal timesteps. They also find that both score-based generative models (Song & Ermon, 2019) and that DDPMs can be formulated by stochastic differential equations with different discretization. Some variants of DDPMs are recently introduced, including the variational diffusion model (VDM) Kingma et al. (2021) that also learns the forward process, and consistency model (Song et al., 2023) and rectified flow model (Liu et al., 2022b), both of which aim to learn to generate natural images in one inference step.
@@ 7 Conclusion
8 Acknowledgements
This work is funded by the CALCULUS project (European Research Council Advanced Grant H2020-ERC-2017-ADG 788506), the Flanders AI Research Program and the China Scholarship Council.
@@ Appendix B Figure Details
Figure 3 shows the mean square error (MSE) between the prediction and ground truth at each step in the backward process. Given an image denoted as $x_{0}$, by applying Equation 2 we could obtain a sequence of $x_{t},t=1,2,\cdots,T-1$. Taking each $x_{t}$ and the paired time step $t$ to run the backward process, we would obtain a sequence of predicted $\hat{x}_{0}^{t},t=1,2,\cdots,T-1$. Ideally, we would expect all these predicted $\hat{x}_{0}^{t},t=1,2,\cdots,T-1$, to be exactly equal to the ground truth $x_{0}$, as they are generated using the given $x_{0}$. However, this is not the case in practice. In our experiments, we found that only when $t<t_{s}$ (around 650 steps in our experiment using DDIM) we could obtain the original $x_{0}$ by running the backward process with paired $(x_{t},t)$. For $t>t_{s}$, the image created using $(x_{t},t)$ differs from the original $x_{0}$. This observation also reveals that the image generation process of diffusion models basically contains two stages. In the first stage, the model moves the Gaussian distribution towards the image distribution and no modes are presented at this stage, which means we can not know which images will be generated. In the second stage, the prediction shows that clear patterns and modes are presented. We can predict which images will be generated following the backward process. This observation led us to divide the error computation into two stages. The full explanation, including the equations, is shown in Figure 8
@@ Appendix C Additional Experimental Setup
Following DDIM (Song et al., 2021), we use two types of time step selection procedures, namely uniform and quadratic. For comparing different sampling methods, the architecture of our models follows the one of DDPM (Ho et al., 2020). For all datasets, we use the same pretrained models for evaluating the sampling methods. For CIFAR-10 and LSUN-bedroom, we obtain the checkpoints from the original DDPM implementation; for CelebA, we adopt the DDIM checkpoints. To carry out the comparison between ADM (Dhariwal & Nichol, 2021), ADM-IP (Ning et al., 2023) and our method, we choose the ADM architecture and the checkpoints from ADM-IP.
@@ Appendix D Error Analysis
We provide here more examples of prediction errors on training samples for different sampling steps for the different datasets. Figure 9 presents the prediction errors obtained in the CelebA dataset, which show similar patterns as those of the CIFAR-10 dataset as dipicted in Figure 3.
@@ Appendix F Sampling Time Comparison
We report a detailed comparison of sampling time using different sampling methods in Table 3 and 4.
@@ Appendix G Case Study
The proposed novel sampling method is motivated by the theoretical and empirical analysis of the exposure bias problem in DDPM. Experimental results show that our sampling method can effectively improve the quality of the generated images compared to images generated with the original DDIM/DDPM models quantatively measured with the FID score. We present the comparison of the generation chain of TS-DDIM (our method) and DDIM in Figure 11. It can be seen that TS-DDIM shifted most of the time steps before reaching the cutoff value, which yields a much better generation quality of the image of a horse. Additional qualitative examples will be presented in the Appendix I.
@@
Figure 12: Comparison of timestep trajectory for TS-DDIM and DDIM on CIFAR-10 with DDPM as backbone using 10 sampling steps and uniform time selection. We use the red dashed line and blue line to present the time step trajectories for the original DDIM and our TS-DDIM, respectively. To improve visibility, we also zoom in for the time steps between 300 to 600.
@@ Appendix H Qualitative Examples
In this section we present the example images generated using different sampling methods with uniform time selection procedure for varying sampling time steps. Generated examples can be found in Figure 13,  14,  15,  16,  17 and  18. It can be seen that we can generate images with good quality for less than 10 sampling steps.
@@ Appendix I Additional Results
We present more results obtained with the LSUN-bedroom dataset (Yu et al., 2016) in Table 5. Limited by computational resources, we do not properly tune the parameters, i.e., window size and cutoff values, for the LSUN-bedroom. We leave the exploration of a more efficient tuning strategy on high-resolution images for future work.
@@
Table 6: Comparison of TS-DDIM with Analytic-DPM on CIFAR-10 (32×32) and CelebA (64×64)
@@
We further integrate our method into the DPM-solver (Lu et al., 2022) and the DEIS (Zhang & Chen, 2023) to showcase its versatility. The results are presented in Table 9. Our model can still improve the performance of both DPM-solver and DEIS samplers. However, in comparison to DDIM, our method again provides substantial improvements although the improvements for these samplers are a bit smaller than for the other tested samplers.This could be attributed to the fact that both the DPM-solver and DEIS utilize the particular structure of the semi-linear ODE, which already largely reduces the error of estimating $x_{t}$.
@@
Finally, in Table 10, we report the performance of our method performed on text-to-image generation on MSCOCO val2017.
@@
In this section, we prove Theorem 3.1.
@@
$=\frac{1}{2}(d\log(1-\overline{\alpha}_{t_{s}})+\frac{1}{1-\overline{\alpha}_{t_{s}}}Tr(\hat{\Sigma}_{t-1})+\frac{1}{1-\overline{\alpha}_{t_{s}}}||\hat{\mu}_{t-1}-\mu_{t_{s}}||^{2})+C$
@@
$=\frac{1}{2}(d\log(1-\overline{\alpha}_{t_{s}})+\frac{1}{1-\overline{\alpha}_{t_{s}}}Tr(\hat{\Sigma}_{t-1})+\frac{1}{1-\overline{\alpha}_{t_{s}}}||(\sqrt{\overline{\alpha}_{t-1}}-\sqrt{\overline{\alpha}_{t_{s}}})x_{0}+e||^{2})+C$
@@
$\mathcal{D}_{KL}(p(\hat{x}_{t-1}|\hat{x}_{t})||q(x_{t_{s}}|x_{0}))\approx\frac{1}{2}(d\log(1-\overline{\alpha}_{t_{s}})+\frac{1}{1-\overline{\alpha}_{t_{s}}}Tr(\hat{\Sigma}_{t-1})+\frac{1}{1-\overline{\alpha}_{t_{s}}}||e||^{2})+C$ (12)
@@
$\sigma_{t_{s}}=\frac{1}{d}Tr(\hat{\Sigma}_{t-1})+\frac{1}{d}||e||^{2}$ (13)
@@
In the above derivation, we assume that when $t$ is large Equation 21 holds. This assumption corresponds to the cutoff mechanism in our proposed algorithm, where we stop conducting time shift when $t$ is small, as the assumption does not hold and we are not able to estimate the $\hat{\Sigma}_{\hat{x}_{t-1}}$ in Equation 13.